Solve the differential equation of the following homogeneous equation.

(x^2 - xy + y^2)dx - xydy = 0
1
Expert's answer
2021-09-17T04:12:50-0400

"(x^2 -xy +y^2)dx -xydy =0 \\
Divide \ by \ x^2 \\

(1-\frac{y}{x} +\frac{y^2}{x^2}) dx - \frac{y}{x} dy = 0 \\
Let \ \frac{y}{x} = v \\
y=vx \\

(1-v+v^2) dx - v dy = 0 \\ 
dy = (vdx+xdv) \\
(1-v+v^2) dx - v (v dx+x dv) = 0 \\ 
(1-v+v^2-v^2) dx - vx dv = 0 \\
(1-v) dx - vx dv = 0 \\
(1-v) dx = vx dv \\

Divide \ both \ sides \ by \ (1-v) x\\ 
\frac{dx}{x} = \frac{v}{(1-v)} dv \\ 

Integrate \ both \ sides\ , we \ get: \\ 
\frac{v}{(1-v)} = \frac{1}{(1-v)} - 1 \\

∫ \frac{dx}{x} = ∫ \frac{dv}{(1-v)} - ∫ dv \\

Substitute \ v=\frac{y}{x}\\ 

ln(x) + C_1 = -ln(1-v) - v \\ 
ln(x) + C_1 = - ln[\frac{(x-y)}{x}] - \frac{y}{x}\\ 
ln[\frac{(x-y)}{x}] + \frac{y}{x} = - ln(x) + C"